OpenStudy (brekathleen):

if f(x)= 2x^2 find f(5) answers 102 52 22 -1.5 please and thank you (:!

OpenStudy (shamim):

Given f(x)=2*x^2 So f(3)=2*3^2

OpenStudy (shamim):

Just substitute x=3

OpenStudy (shamim):

For f(5)=?

OpenStudy (shamim):

Just put x=5 in ur given equation